5. The following information pertains to the Moline Facility for the month of May (all materials are added at the beginn...

5.

The following information pertains to the Moline Facility for the month of May (all materials are added at the beginning of the process):

Units Material Costs
Beginning work in process 80,000 $ 29,000
Started in May 280,000 97,800
Units completed 260,000
Ending work in process 100,000

Required:

Compute the cost per equivalent unit for materials using the weighted-average method. (Round "Cost per equivalent unit" answer to 2 decimal places.)
Answer #1

Equivalent unit of material = Unit transferred out+Ending WIP = 260000+100000 = 360000

Cost per equivalent unit of material = (29000+97800)/360000 = 0.35